OpenStudy (anonymous):

f(y)=y+2/y^2-2y domain of rational function PLEASE HELP

OpenStudy (anonymous):

\[f(y)=\frac{ y+2 }{ y^2-2y }\]

OpenStudy (mathstudent55):

What values can y have? Since there is an expression with y in the denominator, that expression cannot equal zero. Set the expression equal to zero, then solve the equation. Thise are the values y cannot have. The domain is all the values y can have. y^2 - 2y = 0 y(y - 2) = 0 y = 0 or y -2 = 0 y = 0 or y = 2 Therefore, the domain is all real numbers except y=0, y=2

OpenStudy (anonymous):

can you help me with this one I am looking for the range on this problem, I got the domain but can not understand the range\[f(x)=\sqrt{x-3}\]